# CCD-Bench: Probing Cultural Conflict in Large Language Model Decision-Making

### v1

@@ Introduction
To assess similarities in cultural selection patterns and their relation to developers’ locations, we compute symmetrized Kullback-Leibler (KL) divergence between each pair of models’ cluster distributions, yielding a $\times17$ distance matrix fed into average-linkage hierarchical clustering. Developer geographies are from public sources: US (Anthropic, Google, Meta, Microsoft, OpenAI, Perplexity, xAI), China (DeepSeek, Qwen), Canada (Cohere), France (Mistral), and Israel (AI21). Figure 4 shows the KL-divergence heatmap and dendrogram.
@@
The dendrogram in Figure 4 reveals tight intra-family groupings at low divergence levels (e.g., Meta’s Llama-3.3 and Llama-4 at 0.0033, Anthropic’s Claude-3.7 and Claude-4 at 0.0398, Google’s Gemini-2.0 and Gemini-2.5 at 0.0103, OpenAI’s GPT-4o and GPT-4.1 at 0.0200, Microsoft’s Phi-4 and GPT-4.1 at 0.0032 via linkage). Broader clusters emerge around 0.01–0.05: one subgroup includes Google, OpenAI, xAI, Cohere, and Qwen (divergences <<0.03 within); another features Anthropic, Meta, Perplexity, Mistral, and DeepSeek (similarly low internal divergences).
@@
Clustering shows weak geographic ties: US models span both subclusters, with corporate lineage (e.g., Meta/Anthropic vs. Google/OpenAI) dominating over location. Non-US models fragment similarly (Chinese: DeepSeek in subcluster 2, Qwen in 1; European-adjacent: Mistral in 2, Cohere in 1; AI21 with Google/Cohere). This implies that alignment strategies and shared corpora influence cultural behavior more than national origin, though Western-centric data limits distinctions. A notable limitation is our exclusive use of English prompts, which may bias outputs toward Anglo-American values and obscure regional nuances (Zhong, Yun, and Sun 2024); future work should utilize multilingual prompts to distinguish between language and cultural effects.
@@ Conclusion
In this work, we introduce CCD-Bench, a novel benchmark specifically designed to assess how contemporary LLMs navigate conflicts among legitimate cultural values. Evaluating 17 LLMs on 2,182 dilemmas, we observed a pronounced cultural trend toward Nordic and Germanic clusters, hallmarks of egalitarian, consensus-oriented reasoning, while perspectives rooted in Eastern Europe and the Middle East & North Africa remained markedly under‐represented. Despite high rates of multi-dimensional rationales, models relied on a narrow set of GLOBE dimensions, while rarely invoking Assertiveness or Gender Egalitarianism (<<3% each), suggesting an aversion to confrontational or gender-focused reasoning and indicating limited true pluralism in reasoning. Ordering biases were negligible due to our Stratified Latin Square design, and model similarities clustered more by developer lineage than geographic provenance.
These results underscore the need for more nuanced approaches to cultural dynamics in LLMs if they are to operate credibly within an increasingly pluralistic world. Future research could extend CCD-Bench to multilingual prompts, incorporate dynamic user interactions, or explore mitigation strategies, such as targeted fine-tuning, to foster more diverse and adaptable decision-making.

### v2

@@ Introduction
To assess similarities in cultural selection patterns and their relation to developers’ locations, we compute symmetrized Kullback-Leibler (KL) divergence between each pair of models’ cluster distributions, yielding a $\times17$ distance matrix fed into average-linkage hierarchical clustering. Developer geographies are from public sources. Figure 4 shows the KL-divergence heatmap and dendrogram.
@@
The dendrogram in Figure 4 reveals tight intra-family groupings at low divergence levels (e.g., Meta’s Llama-3.3 and Llama-4 at 0.0033, Anthropic’s Claude-3.7 and Claude-4 at 0.0398, Google’s Gemini-2.0 and Gemini-2.5 at 0.0103, OpenAI’s GPT-4o at 0.0200, Microsoft’s Phi-4 and GPT-4.1 at 0.0032 via linkage). Broader clusters emerge around 0.01–0.05: one subgroup includes Google, OpenAI, xAI, Cohere, and Qwen; another features Anthropic, Meta, Perplexity, Mistral, and DeepSeek.
@@
Clustering shows weak geographic ties: US models span both subclusters, with corporate lineage (e.g., Meta/Anthropic vs. Google/OpenAI) dominating over location. Non-US models fragment similarly (Chinese: DeepSeek in subcluster 2, Qwen in 1; European-adjacent: Mistral in 2, Cohere in 1; AI21 with Google/Cohere). This implies that alignment strategies and shared corpora influence cultural behavior more than national origin, though Western-centric data limits distinctions.
Discussion
Our findings reveal a clear cultural bias in contemporary LLMs: when presented with plural legitimate options in a decision-making task, models tend to gravitate toward Nordic/Germanic clusters. The associated rationals are strikingly uniform: models rely on Future Orientation and Performance Orientation with regularity, while rarely grounding decisions in Assertiveness or Gender Egalitarianism (each $3\%$). This does not merely imply that models have bias; rather, the very style of advice being normalised in highly used models embodies a conciliatory, progress-centred, achievement-first worldview, which is systematically misaligned with contexts that legitimate hierarchy or where an explicit appeal to gender fairness is the appropriate move. Therefore, occasionally, a default to consensus-seeking performance talk can result in advice that comes across as naive or disrespectful. Models should instead surface plural options, making trade-offs explicit and labeling the value frames that motivate each option, allowing users to decide which options are valid given their local norms.
A notable limitation is our exclusive use of English prompts, which may bias outputs toward Anglo-American values and obscure regional nuances (zhong_cultural_2024); future work should utilize multilingual prompts to distinguish between language and cultural effects. If LLMs are to mediate everyday social decision-making, alignment must treat conflicts among legitimate values as a design target, so that a model helping someone write to their supervisor or address a family dispute can offer guidance that is not simply good, but good for their values.
@@ Conclusion
In this work, we introduce CCD-Bench, a novel benchmark specifically designed to assess how contemporary LLMs navigate conflicts among legitimate cultural values. Evaluating 17 LLMs on 2,182 dilemmas, we observed a pronounced cultural trend toward Nordic and Germanic clusters. Despite high rates of multi-dimensional rationales, models relied on a narrow set of GLOBE dimensions, while rarely invoking Assertiveness or Gender Egalitarianism (<<3% each), suggesting an aversion to confrontational or gender-focused reasoning and indicating limited true pluralism in reasoning. Our results underscore the need for more nuanced approaches to cultural dynamics in LLMs if they are to operate credibly within an increasingly pluralistic world.
@@ Acknowledgments
This work is supported in part by the NYUAD Center for Interdisciplinary Data Science & AI, funded by Tamkeen under the NYUAD Research Institute Award CG016.